Consider the following data set: 20, 30, 38, 22, 28, 50, 60. What is the IQR of this data set?
Answer #1

Soly tion 20, 30 38 22, 28, Sor Go Deta set I @R=03- 6 Formula First quarhile (ntLyth obserlation valye Third guarhle 03= (3(2nd obsn valuc 11 二 22 O1 22 C03(30nH) )h obsh value 4 6th obsh value O3ニ50 HOR= 03- 0 So-22 HOR 28